OpenStudy (anonymous):

solve by factoring: 2x^2-13x+20=0

OpenStudy (anonymous):

2x^2 -8x -5x +20 2x(x-4)-5(x-4) (2x-5)(x-4) x= 4 or x= 5/2

OpenStudy (anonymous):

(2x-5)(x-4)

OpenStudy (anonymous):

where did you get the 8x and 5x from?

OpenStudy (anonymous):

factor by grouping: what two numbers multiply to (a*c) or (2*20=40) and add to (b) or (-13)? -8 and -5 ax^2+bx+c

OpenStudy (anonymous):

oh okay I get it thanks a lot :)